OpenStudy (anonymous):

whaats x+y=5? it says looking for y terms in terms of x?

OpenStudy (anonymous):

when it says find y in terms of x it means solve for y.

OpenStudy (anonymous):

how do you find it?

OpenStudy (anonymous):

you need to get y by itself so you subtract x from both sides.

OpenStudy (anonymous):

so then it would be y=5-x?

OpenStudy (anonymous):

yep, but you want to turn that into y=-x+5 by switching -x and 5 around, because x usually comes first in an answer. You got it right, though.

OpenStudy (anonymous):

okay so whats after that ? :)

OpenStudy (anonymous):

nothing. That is your answer. You know y is equal to -x+5, so you solved for y in terms of x.

OpenStudy (anonymous):

you turn it into y = -x + 5 and your done if you solving for y

OpenStudy (anonymous):

okay thanks:)